# Understanding Pan-Sharpening via Generalized Inverse

### v2

@@ I Introduction
Deep learning (DL) techniques have become prevalent in MS Pan-sharpening and HS image fusion. The first DL-based Pan-sharpening approach, [18], adapted an autoencoder inspired by sparse denoising. In 2016,[19] introduced the first fully convolutional neural network (CNN) for Pan-sharpening, termed the Pan-sharpening neural network (PNN), which used a three-layer architecture inspired by a super-resolution CNN. Also in 2016, [20] proposed a compressed sensing pansharpening method using the Gram-Schmidt transform and a pre-existing super-resolution CNN. These initial studies spurred significant research interest in DL-based Pan-sharpening, leading to numerous publications. Deep learning methods[21],[22],[23], [24],[25],[26],[27],[28],[29],[30],[31],[32],[33] incorporate more prior knowledge for Pan-sharpening tasks. Consequently, current research focuses on developing novel network architectures and pre-processing techniques to enhance generalization.
@@
  • Integration of generalized inverse theory with CS and MRA methods to obtain two general solution spaces, bridging the gap between mathematical theory and pansharpening techniques.
  • Derivation of the precise correspondence between GSA methods and the generalized inverse expression of CS methods.
  • Improvement in the estimation of the spatial and spectral response matrices using downsampling enhancement.
  • Embedding of prior knowledge in the generalized inverse of the spectral response matrix for sharper reconstructed images.
  • Incorporation of diffusion model prior knowledge for refined pan-sharpening results.
Since we have discovered that the two solutions derived using generalized inverse theory correspond to the CS and MRA methods, respectively, with a deeper understanding of the generalized inverse, more pan-sharpening methods are likely to be discovered. Under the guidance of the derived solution spaces and deep learning generative priors, traditional pan-sharpening results can be further refined.
@@ II-H1 Component substitution form analysis
If $\mathbf{X}_{recover}=\mathbf{V}\mathbf{Z}+(\mathbf{Y}-\mathbf{V}\mathbf{Z}\mathbf{A})\mathbf{W}$, then $\mathbf{X}_{recover}\mathbf{A}-\mathbf{Y}=(\mathbf{I}-\mathbf{W}\mathbf{A})(\mathbf{V}\mathbf{Z}\mathbf{A}-\mathbf{Y}).$
@@ III-E Comparative experiments: performance comparison on WV3 New York
In the experiment, bilinear down-sampling was selected as the spatial down-sampling matrix for calculating the spectral RMSE. The spectral weighting matirx was computed using our proposed method.
@@ III-E1 Quantitative results
TABLE I: Comparison experiments on WV3 New York Dataset
@@
EXP[42] 4.5452 109.3282
BT-H[13] 71.6169 13.2523
BDSD-PC[15] 114.9423 74.0747
C-GSA[12] 75.4264 22.7292
SR-D[16] 100.4333 104.7271
MTF-GLP-HPM-R[5] 74.2957 62.0148
MTF-GLP-FS[7] 71.2183 56.1487
TV[17] 97.7862 59.4001
PanNet[49] 83.1296 70.5015
DRPNN[25] 83.2450 44.5463
MSDCNN[26] 84.7197 48.4944
BDPN[28] 85.6572 43.0467
DiCNN[27] 88.2270 74.2025
PNN[19] 121.9181 71.8810
APNN[29] 79.4560 51.3228
FusionNet[30] 107.6745 120.4564
@@ III-E2 Qualitative results
Figure 2: Comparisons of the different pansharpening methods[42] for WV3 New York dataset.
@@ III-F1 Generation of the simulated data
TABLE II: Chikusei synthetic experiment
@@ VI-A Proofs
$=$ $(\mathbf{I}-\mathbf{W}\mathbf{A})(\mathbf{V}\mathbf{Z}\mathbf{A}-\mathbf{Y}).$

### v3

@@ I Introduction
Deep learning (DL) techniques have become prevalent in MS Pan-sharpening and HS image fusion. The first DL-based Pan-sharpening approach, [18], adapted an autoencoder inspired by sparse denoising. In 2016,[19] introduced the first fully convolutional neural network (CNN) for Pan-sharpening, termed the Pan-sharpening neural network (PNN), which used a three-layer architecture inspired by a super-resolution CNN. Also in 2016, [20] proposed a compressed sensing Pan-sharpening method using the Gram-Schmidt transform and a pre-existing super-resolution CNN. These initial studies spurred significant research interest in DL-based Pan-sharpening, leading to numerous publications. Deep learning methods[21],[22],[23], [24],[25],[26],[27],[28],[29],[30],[31],[32],[33] incorporate more prior knowledge for Pan-sharpening tasks. Consequently, current research focuses on developing novel network architectures and pre-processing techniques to enhance generalization.
@@
  • The integration of generalized inverse theory with CS and MRA methods yields two general solution spaces, bridging the gap between mathematical theory and Pan-sharpening techniques. This integration helps clarify the implicit objectives of different CS and MRA methods that were previously not fully explored. It provides a unified mathematical framework to understand how these two prominent classes of Pan-sharpening methods operate. By establishing this connection, it enables a deeper exploration of the solution spaces for further optimization.
  • We derive the precise correspondence between GSA methods and the generalized inverse representation of CS methods. This correspondence shows that GSA is a special case within the CS framework’s generalized inverse representation, enhancing the understanding of GSA’s mathematical foundation and its relationship with broader CS methods. It allows GSA to be further refined and optimized using potential prior knowledge, such as diffusion priors.
  • We improve the estimation of spatial and spectral response matrices using down-sampling enhancement. Experiments demonstrate that this enhancement results in better quantitative and qualitative outcomes for both CS and MRA methods.
  • Embedding prior knowledge in the range of the generalized inverse of the spectral response matrix helps obtain sharper reconstructed images. Empirically, it is found that when elements of the generalized inverse are close to 1, the image is sharper, while smaller elements cause blurring. By restricting the generalized inverse’s elements to a range around 1 (0.9-1.4), the method avoids blurred features.
  • Incorporating diffusion model prior knowledge into the generalized inverse solution spaces leads to refined Pan-sharpening results. The general solution spaces of the generalized inverse form are substituted into the evidence lower bound of the diffusion model, and the auxiliary matrix is optimized using Score Distillation Sampling. Experiments validate that the diffusion prior significantly improves performance across almost all evaluation measures. This integration leverages the diffusion model’s prior knowledge to further refine solutions within the general solution spaces.
@@ II-D Solution for $\mathbf{A}$ and $\mathbf{B}$
In the subsequent theoretical discussion of error analysis, the use of down-sampling enhancement will induce the equivalence of the CS method and the MRA method. Experiments demonstrate that this enhancement results in better quantitative and qualitative outcomes for both CS and MRA methods.
@@ II-H1 Component substitution form analysis
If $\mathbf{X}_{recover}=\mathbf{V}\mathbf{Z}+(\mathbf{Y}-\mathbf{V}\mathbf{Z}\mathbf{A})\mathbf{W}$, then $\mathbf{X}_{recover}\mathbf{A}-\mathbf{Y}=(\mathbf{V}\mathbf{Z}\mathbf{A}-\mathbf{Y})(\mathbf{I}-\mathbf{W}\mathbf{A}).$
@@ III-E Comparative experiments: performance comparison on WV3 New York
In the experiment, bilinear down-sampling was selected as the spatial down-sampling matrix for calculating the spectral RMSE. The spectral weighting matirx was computed using our proposed method. Here we mainly present the full-resolution experiments to highlight the good consistency of our method.
@@ III-E1 Quantitative results
TABLE I: Full-resolution comparison experiments on WV3 New York Dataset
@@
EXP[42] 4.5452 109.2157
BT-H[13] 71.6169 17.8711
BDSD-PC[15] 114.9423 74.8058
C-GSA[12] 75.4264 26.8394
SR-D[16] 100.4333 107.7014
MTF-GLP-HPM-R[5] 74.2957 63.7501
MTF-GLP-FS[7] 71.2183 58.1442
TV[17] 97.7862 61.4820
PanNet[49] 83.1296 71.4575
DRPNN[25] 83.2450 42.0010
MSDCNN[26] 84.7197 45.3580
BDPN[28] 85.6572 39.7455
DiCNN[27] 88.2270 73.7617
PNN[19] 121.9181 69.4336
APNN[29] 79.4560 52.2568
FusionNet[30] 107.6745 122.9544
@@ III-E2 Qualitative results
Figure 2: Comparisons of the different Pan-sharpening methods[42] for WV3 New York dataset.
@@ III-F1 Generation of the simulated data
TABLE II: Chikusei synthetic experiments
@@ VI-A Proofs
$=$ $(\mathbf{V}\mathbf{Z}\mathbf{A}-\mathbf{Y})(\mathbf{I}-\mathbf{W}\mathbf{A}).$
@@
VI-B Detailed dataset descriptions for Chikusei dataset and PAirMax dataset
Detailed dataset description for the Chikusei dataset.[43] This airborne hyperspectral dataset, acquired by a Headwall Hyperspec-VNIR-C sensor over Chikusei, Ibaraki, Japan, on July 29, 2014 (9:56-10:53 UTC+9), features 128 bands covering 363 nm to 1018 nm. The scene (center: 36.294946°N, 140.008380°E) covers agricultural and urban areas with dimensions of 2517x2335 pixels and a 2.5 m ground sampling distance. Field surveys and visual interpretation of high-resolution Canon EOS 5D Mark II images collected alongside the flight provided ground truth labels for 19 classes(not used). The dataset and ground truth are distributed in ENVI and MATLAB formats for scientific use.
Detailed dataset description for the PAirMax dataset.[40] The PAirMax dataset consists of 14 scenes selected as representative examples of the heterogeneity encountered in Pan-sharpening applications. The dataset includes 14 Full-Resolution (FR) datasets. 14 Reduced-Resolution (RR) datasets were generated from these FR test cases. Most images depict urban areas, presenting significant pansharpening challenges such as accurately rendering high-contrast features (e.g., building edges), sub-resolution details, and avoiding spectral smearing across adjacent regions with differing spectral properties. Some scenes feature natural land cover types, including vegetation (agricultural fields, meadows, forests), and water bodies. Vegetated areas pose challenges due to the distinct reflectance difference between visible and NIR wavelengths (caused by chlorophyll) and textured patterns with sub-pixel size, potentially leading to spectral and spatial distortions. Water regions are included to assess spectral reconstruction, particularly relevant for sensors with bathymetry-optimized bands (e.g., WorldView-2/3), and often feature sharp edges against urban embankments. The scenes were acquired across different seasons, introducing variations in illumination conditions, such as dimmer intensities and elongated shadows in winter imagery. All PAN-MS bundles exhibit a spatial resolution ratio of 4:1 (PAN has 16 times more pixels than each MS band). MS bands cover the Visible and Near-Infrared (VNIR) domain, with either four bands (GeoEye-1, WorldView-2, SPOT-7) or eight bands (WorldView-2/3). The 14 images in PAirMax were derived from original PAN+MS bundles acquired under clear-sky conditions with negligible cloud cover (especially within the selected cropped areas).
VI-C Synthetic PAN and HS images in synthetic experiments.
(a) Generated pan
(b) RGB
(c) OAP
VI-D Down-sampling and up-sampling parameters illustration
In the comparative experiment, we use bilinear down-sampling and bilinear up-sampling. In the synthetic experiment, we use mean down-sampling to create the data and use bilinear down-sampling and bilinear up-sampling for calculation of our method. In ablation study and diffusion-related experiment, we use mean down-sampling and nearest neighborhood up-sampling.
VI-E Reduced-resolution comparative experiments
In the reduced-resolution comparative experiments, our method does not perform outstandingly because it has not been refined by searching the general solution space with the diffusion prior.
TABLE V: Reduced-resolution comparative experiments on WV3 New York Dataset
Method Q2n($\uparrow$) SAM($\downarrow$) ERGAS($\downarrow$)
GT 1.0000 0.0000 0.0000
EXP 0.6555 7.2344 8.2676
BT-H 0.9261 6.4689 3.9948
BDSD-PC 0.9334 6.8699 3.9157
C-GSA 0.9235 6.7104 4.0751
SR-D 0.9098 6.6782 4.4543
MTF-GLP-HPM-R 0.9245 7.0266 4.0949
MTF-GLP-FS 0.9246 6.7789 4.0691
TV 0.9294 6.6165 4.0824
PanNet 0.9217 6.9522 4.3506
DRPNN 0.9224 7.4331 4.2887
MSDCNN 0.9093 7.5705 4.4867
BDPN 0.9198 7.7437 4.4823
DiCNN 0.8558 8.1309 5.6302
PNN 0.8830 12.6763 6.8061
APNN 0.9119 7.6778 4.5471
FusionNet 0.8534 8.4185 6.1608
PCS/PMRA 0.8703 9.9481 5.6316